calculate the bond of return for a bond with face value of USD 1000 that pays an 8% coupon for USD 1070 and sell it in a year's time for USD 1190
Answer #1

Ans:

Rate of return = ( Gain for holding the bond + Interest ) / Purchase price

= ( Selling price - purchase price ) + Interest / Purchase price

= ( 1190 - 1070 ) + ( 1000 * 8% ) / 1070

= ( 120 + 80 ) / 1070

= 200 / 1070

= 0.1869

= 18.69 %